# Edge AI: A Taxonomy, Systematic Review and Future Directions

### v1

@@ Abstract
Edge Artificial Intelligence (AI) incorporates a network of interconnected systems and devices that receive, cache, process, and analyse data in close communication with the location where the data is captured with AI technology. Recent advancements in AI efficiency, the widespread use of Internet of Things (IoT) devices, and the emergence of edge computing have unlocked the enormous scope of Edge AI. The goal of Edge AI is to optimize data processing efficiency and velocity while ensuring data confidentiality and integrity. Despite being a relatively new field of research, spanning from 2014 to the present, it has shown significant and rapid development over the last five years. In this article, we present a systematic literature review for Edge AI to discuss the existing research, recent advancements, and future research directions. We created a collaborative edge AI learning system for cloud and edge computing analysis, including an in-depth study of the architectures that facilitate this mechanism. The taxonomy for Edge AI facilitates the classification and configuration of Edge AI systems while also examining its potential influence across many fields through compassing infrastructure, cloud computing, fog computing, services, use cases, ML and deep learning, and resource management. This study highlights the significance of Edge AI in processing real-time data at the edge of the network. Additionally, it emphasizes the research challenges encountered by Edge AI systems, including constraints on resources, vulnerabilities to security threats, and problems with scalability. Finally, this study highlights the potential future research directions that aim to address the current limitations of Edge AI by providing innovative solutions.
@@ I Introduction
Recent advancements in Artificial Intelligence’s (AI) effectiveness, the adoption of Internet of Things (IoT) devices, and edge computing capabilities are coming together to unleash the potential of edge artificial intelligence (Edge AI) [1]. Numerous analysts and businesses are conversing about and executing edge computing, which delineates its origins to the 1990s when edge servers positioned near customers were used to serve web and video content over content delivery networks [2]. In this edge AI, Edge computing is a paradigm transformation that brings data storage and processing closer to the data source, improving response times and reducing bandwidth usage. Unlike traditional cloud computing, where centralized data centres process data, edge computing processes data at the network’s edge [3]. This proximity reduces latency, enhances real-time data processing capabilities, and supports the expansion of IoT devices and services[4]. The primary advantages of edge computing include improved agility of services, low latency, enhanced coherence, and the elimination of a single point of failure, making it highly relevant for applications in smart cities, self-sufficient vehicles, and industrial automation [5]. By distributing computing resources geographically, edge computing ensures that data processing takes place near the data source, satisfying the need for analytics and decision-making in real time.
@@ I-A Edge Computing for AI
Furthermore, AI-based technologies play a vital role in addressing Quality of Service (QoS)-aware scheduling and resource allocation challenges in edge environments, ensuring quality of service and user experience. Edge AI enables the deployment of AI as a Service (AIaaS) with configurable model complexity and data quality, enhancing performance and reducing costs [12, 13]. This innovative approach supports smart security applications by leveraging AI capabilities at the edge and enhancing security measures for distributed systems. Edge intelligence, a promising technology, empowers real-time applications by moving computing from cloud servers to IoT edge devices, creating intelligent enterprises with vast possibilities [14]. The utilization of AI at the edge, instead of centralized locations, unlocks the potential of AI with IoT devices and edge computing, deploying AI algorithms on resource-constrained edge devices for various applications like autonomous vehicles, healthcare, and surveillance.
@@ I-B Motivation and Contributions
The main contributions of this paper are:
  • We conducted a systematic review that provided a thorough examination of edge AI research based on many applications, highlighting current trends and possible directions for the future.
@@ I-C Article Organization
Section II discusses a related survey focusing on different applications in terms of algorithms, optimization techniques, security, and privacy concerns integrated with Edge AI. Section III covers the background and present present situation of Edge AI. Section IV details the methodology adopted for the review. Section V outlines a taxonomy encompassing infrastructure, cloud computing, fog computing, services, use cases, machine learning and deep learning, and resource management. Section VI compares existing Edge AI implementations based on taxonomy. Section VII presents an analysis and the results obtained, and the future research directions are discussed in Section VIII. Finally, Section IX summarizes the survey.
II Related Surveys
@@
In this section, we discuss various studies, which are about the different applications consisting of smart cities, smart manufacturing, autonomous vehicles, the Internet of Vehicles (IoV), industrial automation, and healthcare monitoring systems. These are highlighted when edge computing meets AI and AI for edge computing. There are also considerations of traditional ML, computation offloading optimization, and concerns related to privacy and security, reflecting a comprehensive analysis of the challenges and strategies integrating AI and edge computing.
II-A Smart Cities
@@ II-B Smart Manufacturing
In intelligent manufacturing, combining AI with Edge Computing significantly enhances efficiency, decision-making, and security, effectively addresses challenges, and improves data utilization. This integration streamlines operations, enables predictive maintenance, and supports autonomous decisions. The following papers are discussed to describe these advancements and subsequent challenges: In 2022, authors [29] explored the scope of AI and its application in India’s intelligent manufacturing industry, concentrating on the technology’s current state, constraints, and recommendations for resolving issues. In [30], authors have talked about how ML and AI may boost productivity, sustainability, and manufacturing efficiency. However, there are a number of difficulties with implementing AI in manufacturing, including problems with infrastructure and human resources, security threats, trust, and data management and acquisition. Further, the authors [31] suggested a new mode called ”AI-Mfg-Ops” (AI-enabled Manufacturing Operations) with a supporting software-defined framework is proposed as part of an open evolutionary architecture of the intelligent cloud manufacturing system. This mode can facilitate quick operation and upgrades of cloud manufacturing systems with intelligent assessment, analysis, planning, and execution in a closed loop. In 2021, authors [32] addressed the job shop scheduling challenge in the intelligent factory process while using a Deep Q-network (DQN). The suggested framework is contrasted and examined with other frameworks from the standpoint of offering an intelligent factory service. Furthermore, the authors [33] look at how it tends to integrate several productivity factors, such as big data analytics, Automation, and Operations Information, which connect machines via open platforms, resulting in real-time reactions to boost efficiency across the supply chain. Moreover, the authors[34] developed a service-oriented information model to standardized describe the functional characteristics and related operational data of heterogeneous manufacturing resources; additionally, a message middleware-based real-time transmission and integration method for high-volume operational field and sensor data is suggested in order to achieve the efficient distribution of related data and remote monitoring of distributed manufacturing resources.Fianlly, the authors [35] discusses the possible advantages and difficulties of a federated learning architecture based on data gathered from 5G MSPs to enable predictive maintenance (PM) in industrial settings.
II-C Autonomous Vehicles and the Internet of Vehicles (IoV)
@@
In the context of autonomous and IoT-enabled vehicles, advancements in control and task optimization are propelled by AI integration and Edge Computing (EC). This synergy supports real-time decision-making, highlighting the importance of AI and EC in addressing challenges like real-time processing and security and optimizing communication and privacy within the IoV. The application of DL and Reinforcement Learning with Multiple Agents (MARL) Underscores the need for efficient solutions in autonomous driving technologies. In these scenarios, the following papers are described: In 2023, authors [36] offers a thorough technical overview of the most recent studies conducted in the areas of lateral, longitudinal, and integrated control strategies for self-driving cars. They also examine a variety of strategies and tactics used to attain accurate steering control while taking longitudinal factors into account. [37] discusses key technologies, applications, solutions, and problems related to integrating Mobile Edge Computing (MEC) and ML in the Internet of UAVs and are covered in-depth by the author’s thorough review. Further, authors [38] examined the most recent research on vehicular data offloading from the standpoint of communication, focusing on vehicle-to-vehicle (V2V), vehicle-to-roadside infrastructure (V2I), and vehicle-to-everything (V2X). The study also identified unresolved research issues in this area and forecasted future directions in the field. Furthermore, the authors [39] suggested a multi-access edge computing (MEC) framework to facilitate the cooperation of digital twins (DTs) into wireless networks and connected cars (CVs) to reduce the unreliability of long-distance communication between edge servers and CVs. Moreover, researchers [40] outlines a DL and edge computing-based vehicle intelligent control system that encourages the broad advancement of automation and intelligent technology. The findings show that the distance between the target and experimental vehicles is extremely close to the anticipated safe space. In [41], authors suggested a safe edge intelligence that combined the advantages of blockchain, local differential privacy (LDP), and federated learning (FL) for automotive networks. The authors in [42] focuses on a fast task execution technique in heterogeneous IoT applications that is powered by AI. This technique reduces decision latency by considering various system parameters, including the task’s execution deadline, the device’s battery level, the channel conditions between mobile devices and edge servers, and the capacity of the edge servers. Finally, the authors [43] combined edge computing and the Web of Things, compared their functions, and showed how edge computing improves the efficiency of real-time IoT applications by focusing on transmission, storage, and computation elements.
II-D Industrial Automation
@@
In industrial automation, several papers discuss revolutionary approaches to enhancing productivity by integrating AI, edge computing, robotics, and data analytics. The relevant papers look over the utilization of the Industrial Cyber Intelligent Control Operating System (ICICICOS), a cloud-edge computing-based system, for AI and industrial automation. It focuses on proposing AI with industrial processes at the edge. It emphasizes strategies for optimizing ML methods, deploying AI models on resource-constrained devices, and addressing security concerns through secure AI microservices at the edge. The relevant papers are described as follows: In 2022, the authors [11] presented a flexible working mechanism by permitting the combined design of data quality ratios (DQRs) and model complexity ratios (MCRs) for the AI tasks and suggested a configurable model deployment architecture for edge AIaaS. Further, the authors [12] provide a systematic overview of methods for addressing the paucity of training data for different kinds of data, and a methodology for addressing data scarcity in cellular networks is suggested. In [14], the authors explore the privacy-enhancing solutions that are now in position, including the technologies, specifications, and process solutions to mitigate these risks. It also looks at privacy threats at various stages of the AI life cycle. Further, the authors [16] offer a thorough overview of edge intelligence and lightweight ML support for upcoming services and applications. The researchers have supplied a thorough analysis of cutting-edge intelligence applications, lightweight ML techniques, and their support for upcoming services and applications. In [44], authors provide a thorough review of AI/ML-based IDS/MDSs and set baseline measurements pertinent to networked autonomous systems, emphasizing the gaps and assessment metrics in the existing research. In [6], the authors wrapped up a thorough analysis of edge computing, covering both the shift to edge AI and related paradigms. Additionally, the history of every alternative put out for edge computing implementation, as well as the Edge AI strategy for putting AI models and algorithms on edge devices, were investigated.
II-E Smart Healthcare

### v2

@@ Abstract
Edge Artificial Intelligence (AI) incorporates a network of interconnected systems and devices that receive, cache, process, and analyze data in close communication with the location where the data is captured with AI technology. Recent advancements in AI efficiency, the widespread use of Internet of Things (IoT) devices, and the emergence of edge computing have unlocked the enormous scope of Edge AI. Edge AI aims to optimize data processing efficiency and velocity while ensuring data confidentiality and integrity. Despite being a relatively new field of research from 2014 to the present, it has shown significant and rapid development over the last five years. This article presents a systematic literature review for Edge AI to discuss the existing research, recent advancements, and future research directions. We created a collaborative edge AI learning system for cloud and edge computing analysis, including an in-depth study of the architectures that facilitate this mechanism. The taxonomy for Edge AI facilitates the classification and configuration of Edge AI systems while examining its potential influence across many fields through compassing infrastructure, cloud computing, fog computing, services, use cases, ML and deep learning, and resource management. This study highlights the significance of Edge AI in processing real-time data at the edge of the network. Additionally, it emphasizes the research challenges encountered by Edge AI systems, including constraints on resources, vulnerabilities to security threats, and problems with scalability. Finally, this study highlights the potential future research directions that aim to address the current limitations of Edge AI by providing innovative solutions.
@@ I Introduction
Recent advancements in Artificial Intelligence (AI), the growing adoption of Internet of Things (IoT) devices, and the rise of edge computing are converging to unleash the full potential of edge AI [1]. Numerous analysts and businesses are conversing about and executing edge computing, which delineates its origins to the 1990s when edge servers positioned near customers were used to serve web and video content over content delivery networks [2]. Edge computing is a paradigm transformation in this edge AI that brings data storage and processing closer to the data source, improving response times and reducing bandwidth usage. Unlike traditional cloud computing, where centralized data centers process data, edge computing processes data at the network’s edge [3]. This proximity reduces latency, enhances real-time data processing capabilities, and supports the expansion of IoT devices and services[4]. The primary advantages of edge computing include improved agility of services, low latency, enhanced coherence, and the elimination of a single point of failure, making it highly relevant for applications in smart cities, self-sufficient vehicles, and industrial automation [5]. By distributing resources geographically, edge computing ensures that data processing occurs near the data source, satisfying the need for analytics and decision-making in real-time.
@@ I-A Edge AI
Furthermore, AI-based technologies play a vital role in addressing Quality of Service (QoS)-aware scheduling and resource allocation challenges in edge environments, ensuring QoS and user experience. Edge AI enables the deployment of AI as a Service (AIaaS) with configurable model complexity and data quality, enhancing performance and reducing costs [12, 13]. This innovative approach supports smart security applications by leveraging AI capabilities at the edge and enhancing security measures for distributed systems. Edge intelligence, a promising technology, empowers real-time applications by moving computing from cloud servers to IoT edge devices, creating intelligent enterprises with vast possibilities [14]. The utilization of AI at the edge, instead of centralized locations, unlocks the potential of AI with IoT devices and edge computing, deploying AI algorithms on resource-constrained edge devices for various applications like autonomous vehicles, healthcare, and surveillance.
@@ I-C Article Organization
Section II offer present situation of Edge AI. Section IV details the methodology adopted for the review. Section III discusses a related surveys and studies focusing on different applications in terms of algorithms, optimization techniques, security, and privacy concerns integrated with Edge AI. Section V outlines a taxonomy encompassing infrastructure, cloud computing, fog computing, services, use cases, machine learning and deep learning, and resource management. Section VI compares existing Edge AI implementations based on taxonomy. Section VII presents an analysis and the results obtained, and the future research directions are discussed in Section VIII. Finally, Section IX summarizes the survey.
II Edge AI: Background and Current Status
This section explains some concepts related to background and current status in Edge AI. Subsection II-A explains edge computing and its historical emergence. Subsection II-B provides information on the integration of AI and edge technologies. This section is completed by explaining Edge AI applications and challenges in subsection II-C and subsection II-D, respectively.
II-A Historical Emergence of Edge Computing
The concept of edge computing is a paradigm that brings computing resources closer to the data source, unlike the cloud, which provides services through a remote server [22]. In this way, it is aimed to reduce problems such as unnecessary bandwidth occupation and latency in today’s world where huge amounts of data that need to be processed are produced [23]. To understand the emergence of edge computing, it will be more useful to examine previous paradigms such as cloud and fog computing. Fig. 1 shows the advantages of cloud, fog, and edge computing over each other and their layer arrangement. These concepts are discussed briefly below:
Figure 1: Computing Paradigms and their Objectives.
  • Cloud Computing: It is a paradigm that dates back to the 1970s and refers to the use of common computing resources by users on a server via the Internet [24]. Today, it is offered to users with various service models, especially by large companies such as Microsoft Azure, Google Cloud Platform and IBM Cloud. The advantages of cloud computing are as follows [25]:
    • Scalability, in case the need for computing resources increases (demand fluctuations), cloud computing provides services such as more processing power and storage by scaling the resources. In this way, performance measures such as service-level agreement (SLA) and QoS are ensured.
  • Fog Computing: The concept of fog computing was introduced by Cisco in 2012 [26]. This paradigm recommends moving computing resources closer to the endpoints of the network (such as routers and gateways) to reduce the latency and bandwidth problems that occur in cloud computing. When Fig. 1 is examined, fog computing acts as a layer between the cloud and the edge. The advantages of fog computing are as follows [27]:
II-B Integration of AI with Edge Technology
Developing AI applications have begun to show themselves in many areas. One of these areas is EdgeAI, which is the combination of AI and Edge concepts [29]. EdgeAI is based on the principle of processing data on edge nodes such as mobile devices and IoT instead of processing it on cloud servers [31]. This is achieved by distributing AI algorithms to edge nodes close to the data source, as shown in Figure 2, which shows how data is processed in the EdgeAI concept and how it performs fast and efficient computation. The advantages offered by these two technologies can be listed as follows [6]:
  • Low Latency: In delay-sensitive scenarios such as e-health and autonomous vehicle applications where patients are monitored instantly, millisecond delays are critical [32]. In traditional cloud-based systems, data must be sent between the user and the cloud to be processed in the AI model deployed in the cloud. This process will cause serious delay and unnecessary bandwidth usage [23]. With Edge and AI integration, this problem can be overcome by processing data in real time. Because the data will be processed on the edge node closest to the source where it was produced, it will respond much faster than cloud-based systems.
  • Increased Security and Privacy: In cloud systems, data is sent from the source where it is produced to central servers. This expands the attack surface for hackers in the communication channels and storage areas of sensitive data such as biometric and health data [32]. In EdgeAI systems, since the data is processed and stored locally compared to cloud systems, it can be said that the overall security of the system is higher. Similarly, privacy issues that may occur in the event of theft of sensitive data such as biometric data are reduced [22].
  • Resource Optimization and Scalability: EdgeAI systems consist of heterogeneous devices such as laptops, network routers, and mobile devices with different processing power and storage capabilities. This means that EdgeAI can share the resources of devices in the edge network if external processing power and resources are needed. In addition, balanced load distribution can be achieved by using advanced resource allocation algorithms to optimize resources.
Future Directions and Limitations: Despite the above-mentioned advantages and high potential, EdgeAI also brings with it challenges such as (i) limited processing power of devices in the edge network, (ii) management difficulties due to the heterogeneous structure of the edge network, and (iii) energy constraints due to resource limitations. If future researchers solve these challenges, it is expected that the areas of use of EdgeAI will expand.
Figure 2: Architecture of Edge AI.
II-C Edge AI Applications
Edge AI applications, created by combining the concepts of Edge and AI, provide lower latency and higher security than Cloud-based AI applications. Fig. 3 shows popular applications of Edge AI, which are discussed briefly below:
  • Smart Parking: With the increase in means of transportation, parking has become a big problem, especially in big cities. Edge AI-based solutions with the help of sensors and IoT devices can be used to solve these problems [33].
Figure 3: The Edge AI Applications.
Figure 4: Edge AI Implementation Challenges.
II-D Edge AI Implementation Challenges
EdgeAI, which emerges by combining Edge and AI, brings with it the advantages it offers, but also challenges that are still waiting to be solved. These challenges are shown in Fig. 4. These challenges are discussed briefly below:
  • Energy Efficiency: Edge devices generally consist of homogeneous and heterogeneous devices with low processing and storage capacity. Applications that require Natural Language Processing (NLP) and intensive image processing will cause excessive resource consumption on edge devices [36]. For this reason, new solutions such as special AI chips or task engineering are needed. Therefore, new solutions have emerged, such as dedicated AI chips or task engineering. Examples include Google’s TPU and NVIDIA’s JetsonAI chips, which use low-power algorithms to achieve energy efficiency [37]. Another energy-efficient method is quantization and pruning techniques. These techniques reduce energy consumption by reducing the size of models in neural networks.
  • Maintenance and Updates: Since edge devices consist of devices distributed in different locations, this means more attack targets for hackers [38]. In addition, not all devices in the edge nodes have a homogeneous structure, which means separate system maintenance and updates for each node [39]. Measures such as automatic updating can be taken to solve these problems. In automatic update solutions, system incompatibilities may occur due to the heterogeneous structure of the devices. For this reason, containerization and orchestration solutions come to the fore in terms of service isolation and ease of management [40].
  • Scalability: Since edge devices generally consist of heterogeneous devices, the distribution of a single application to different devices is still a challenge (task scheduling, etc.) [41]. Additionally, it is difficult to synchronize data across all devices. Effective microservice architectures and load-balancing algorithms that prevent a node from being overloaded can be used to solve this problem. Examples include Kubernetes and Istio tools [42]. These tools can easily overcome load balancing and automating service discovery issues, while performance optimization should be considered for scenarios that require low latency.
III Related Studies and Surveys
In this section, we discuss related studies and surveys, as well as our main contributions.
III-A Related Studies
@@
Here, we discuss various studies, which are about the different applications consisting of smart cities, smart manufacturing, autonomous vehicles, the Internet of Vehicles (IoV), industrial automation, and healthcare monitoring systems. These are highlighted when edge computing meets AI and AI for edge computing. There are also considerations of traditional ML, computation offloading optimization, and concerns related to privacy and security, reflecting a comprehensive analysis of the challenges and strategies integrating AI and edge computing.
III-A1 Smart Cities
@@ III-A2 Smart Manufacturing
In intelligent manufacturing, combining AI with Edge Computing significantly enhances efficiency, decision-making, and security, effectively addresses challenges, and improves data utilization. This integration streamlines operations, enables predictive maintenance, and supports autonomous decisions. The following papers are discussed to describe these advancements and subsequent challenges: In 2022, authors [50] explored the scope of AI and its application in India’s intelligent manufacturing industry, concentrating on the technology’s current state, constraints, and recommendations for resolving issues. In [51], authors have discussed how ML and AI may boost productivity, sustainability, and manufacturing efficiency. However, there are several difficulties with implementing AI in manufacturing, including problems with infrastructure and human resources, security threats, trust, and data management and acquisition. Further, the authors [52] suggested a new mode called “AI-Mfg-Ops” (AI-enabled Manufacturing Operations) with a supporting software-defined framework proposed as part of an open evolutionary architecture of the intelligent cloud manufacturing system. This mode can facilitate quick operation and upgrades of cloud manufacturing systems with intelligent assessment, analysis, planning, and execution in a closed loop. In 2021, authors [53] addressed the job shop scheduling challenge in the intelligent factory process while using a Deep Q-network (DQN). The suggested framework is contrasted and examined with other frameworks from the standpoint of offering an intelligent factory service. Furthermore, the authors [54] look at how it tends to integrate several productivity factors, such as big data analytics, Automation, and Operations Information, which connect machines via open platforms, resulting in real-time reactions to boost efficiency across the supply chain. Moreover, the authors[55] developed a service-oriented information model to standardized describe the functional characteristics and related operational data of heterogeneous manufacturing resources; additionally, a message middleware-based real-time transmission and integration method for high-volume operational field and sensor data is suggested to achieve the efficient distribution of related data and remote monitoring of distributed manufacturing resources. Finally, the authors [56] discuss the possible advantages and difficulties of a federated learning architecture based on data gathered from 5G MSPs to enable predictive maintenance (PM) in industrial settings.
III-A3 Autonomous Vehicles and the Internet of Vehicles (IoV)
@@
In the context of autonomous and IoT-enabled vehicles, advancements in control and task optimization are propelled by AI integration and Edge Computing (EC). This synergy supports real-time decision-making, highlighting the importance of AI and EC in addressing challenges like real-time processing and security and optimizing communication and privacy within the IoV. The application of DL and Reinforcement Learning with Multiple Agents (MARL) Underscores the need for efficient solutions in autonomous driving technologies. In these scenarios, the following papers are described: In 2023, authors [57] offer a thorough technical overview of the most recent studies conducted in the areas of lateral, longitudinal, and integrated control strategies for self-driving cars. They also examine a variety of strategies and tactics used to attain accurate steering control while taking longitudinal factors into account. [58] discusses key technologies, applications, solutions, and problems related to integrating Mobile Edge Computing (MEC) and ML in the Internet of UAVs and are covered in-depth by the author’s thorough review. Further, authors [59] examined the most recent research on vehicular data offloading from the standpoint of communication, focusing on vehicle-to-vehicle (V2V), vehicle-to-roadside infrastructure (V2I), and vehicle-to-everything (V2X). The study also identified unresolved research issues in this area and forecasted future directions in the field. Furthermore, the authors [60] suggested a multi-access edge computing (MEC) framework to facilitate the cooperation of digital twins (DTs) into wireless networks and connected cars (CVs) to reduce the unreliability of long-distance communication between edge servers and CVs. Moreover, researchers [61] outline a DL and edge computing-based vehicle intelligent control system that encourages the broad advancement of automation and intelligent technology. The findings show that the distance between the target and experimental vehicles is extremely close to the anticipated safe space. In [62], authors suggested a safe edge intelligence that combined the advantages of blockchain, local differential privacy (LDP), and federated learning (FL) for automotive networks. The authors in [63] focus on a fast task execution technique in heterogeneous IoT applications that are powered by AI. This technique reduces decision latency by considering various system parameters, including the task’s execution deadline, the device’s battery level, the channel conditions between mobile devices and edge servers, and the capacity of the edge servers. Finally, the authors [64] combined edge computing and the Web of Things, compared their functions, and showed how edge computing improves the efficiency of real-time IoT applications by focusing on transmission, storage, and computation elements.
III-A4 Industrial Automation
@@
In industrial automation, several papers discuss revolutionary approaches to enhancing productivity by integrating AI, edge computing, robotics, and data analytics. The relevant papers look over the utilization of the Industrial Cyber Intelligent Control Operating System (ICICICOS), a cloud-edge computing-based system, for AI and industrial automation. It focuses on proposing AI with industrial processes at the edge. It emphasizes strategies for optimizing ML methods, deploying AI models on resource-constrained devices, and addressing security concerns through secure AI microservices at the edge. The relevant papers are described as follows: In 2022, the authors [11] presented a flexible working mechanism by permitting the combined design of data quality ratios (DQRs) and model complexity ratios (MCRs) for the AI tasks and suggested a configurable model deployment architecture for edge AIaaS. Furthermore, the authors [12] provide a systematic overview of methods for addressing the paucity of training data for different kinds of data, and a methodology for addressing data scarcity in cellular networks is suggested. In [14], the authors explore the privacy-enhancing solutions that are now in position, including the technologies, specifications, and process solutions to mitigate these risks. It also looks at privacy threats at various stages of the AI life cycle. Further, the authors [16] offer a thorough overview of edge intelligence and lightweight ML support for upcoming services and applications. The researchers have supplied a thorough analysis of cutting-edge intelligence applications, lightweight ML techniques, and their support for upcoming services and applications. In [65], authors provide a thorough review of AI/ML-based IDS/MDSs and set baseline measurements pertinent to networked autonomous systems, emphasizing the gaps and assessment metrics in the existing research. In [6], the authors wrapped up a thorough analysis of edge computing, covering both the shift to edge AI and related paradigms. Additionally, the history of every alternative put out for edge computing implementation, as well as the Edge AI strategy for putting AI models and algorithms on edge devices, were investigated.
III-A5 Smart Healthcare